OpenStudy (anonymous):

Meryl needs to add enough water to 11 gallons of an 18% detergent solution to make a 12% detergent solution. Which equation can she use to find g, the number of gallons of water she should add?

OpenStudy (anonymous):

Hi !

OpenStudy (anonymous):

Do you have options ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

lookin for equation tho

OpenStudy (anonymous):

Oh ok

OpenStudy (neer2890):

so 18% of 11 gallons = quantity of detergent .18*11=1.96 gallons of detergent. means in 11 gallons quantity of detergent is 1.96 gallons.

OpenStudy (anonymous):

wat was ur equation

OpenStudy (neer2890):

so now we have to find the quantity of mixture(add water but detergent remains 1.96 gallons because we are not adding detergent) in which the %age of detergent is 12%.

OpenStudy (anonymous):

Is it like this x+11 = 12%

OpenStudy (anonymous):

We know 11 = 18% x = 12%

OpenStudy (anonymous):

So how should we put this in to an equation

OpenStudy (neer2890):

so, let x is the quantity of mixture. 12% of x= 1.96 gallons x=16.5

OpenStudy (anonymous):

their no equation?

OpenStudy (neer2890):

first we have to find the answer so that we know what the equation is..!

OpenStudy (neer2890):

but x is the total mixture. in which we know that 11 gallons is of old mixture. so quantity of water added =16.5-11=5.5 gallons

OpenStudy (neer2890):

now lets make an equation of all this.

OpenStudy (neer2890):

let x(new) is the total quantity of mixture in which 12% detergent is there. and x(old) is the old mixture which is 11 gallons. t is the quantity of detergent in the mixture. and g is the number of gallons of water mixed. so g=x(new)-x(old) g= (t/.12)-x(old)

OpenStudy (neer2890):

now here, t=1.96 and x(old)=11 so, g=(1.96/0.12)-11=5.5 gallons